# Holographic Intelligence Surface Assisted Integrated Sensing and Communication

### v1

@@ Abstract
Traditional discrete-array-based systems fail to exploit interactions between closely spaced antennas, resulting in inadequate utilization of the aperture resource. In this paper, we propose a holographic intelligence surface (HIS) assisted integrated sensing and communication (HISAC) system, wherein both the transmitter and receiver are fabricated using a continuous-aperture array. A continuous-discrete transformation of the HIS pattern based on the Fourier transform is proposed, converting the continuous pattern design into a discrete beamforming design. We formulate a joint transmit-receive beamforming optimization problem for the HISAC system, aiming to balance the performance of multi-target sensing while fulfilling the performance requirement of multi-user communication. To solve the non-convex problem with coupled variables, an alternating optimization-based algorithm is proposed to optimize the HISAC transmit-receive beamforming in an alternate manner. Specifically, the transmit beamforming design is solved by decoupling into a series of feasibility-checking sub-problems while the receive beamforming is determined by the Rayleigh quotient-based method. Simulation results demonstrate the superiority of the proposed HISAC system over traditional discrete-array-based ISAC systems, achieving significantly higher sensing performance while guaranteeing predetermined communication performance.
Index Terms:
Holographic intelligence surface, integrated sensing and communication, pattern design strategy, continuous-aperture array
@@ I Introduction
isac system, which envisions simultaneous sensing and communication on a shared waveform/platform, is considered a vital solution for tackling challenges associated with spectral and/or power efficiency in 6G wireless networks [1, 2]. For the ISAC waveform design, former studies advocate a joint design paradigm [3, 4, 5, 6, 7, 8]. Instead of allocating orthogonal temporal/spectral/spatial resources to sensing and communication functionalities [9, 10, 11], joint design achieves ISAC by a single waveform with non-orthogonal resources, which not only enhances spectral efficiency but also yields a flexible performance tradeoff between sensing and communication [3, 4, 5, 6, 7, 8]. In a multi-antenna setting, joint ISAC waveform design is typically achieved through transmit beamforming/precoding design. This entails designing spatial beamformers/precoders judiciously to maximize sensing/communication under performance constraints in communication/sensing, thereby striking a balance between the two functionalities [3, 4, 5]. Furthermore, by leveraging radar receivers, joint transmit-receive beamforming has been proposed in the literature to further enhance ISAC performance [6, 7, 8].
The aforementioned beamforming designs in ISAC systems typically rely on discrete antenna arrays, where the antenna spacing is constrained to half-wavelength. This limits the number of antennas for a given aperture, thereby resulting in insufficient exploitation of array gain for ISAC performance enhancement [12, 13]. In addition, modeling electromagnetic (EM) propagation for the discrete array usually overlooks the mutual interference between two closely spaced antennas, potentially leading to inadequate information acquisition during signal reception [14, 15]. To overcome these limitations, the concept of holographic intelligence surface (HIS), also as known as holographic multiple-input multiple-output (MIMO) [16, 17] and continuous-aperture MIMO [18, 19], has emerged. Unlike the discrete array consisting of multiple patch antennas with half-wavelength spacing, HIS achieves a quasi-continuous-aperture array by deploying massive sub-wavelength tunable elements in a compact space, which yields high array gain with limited size [20]. The advantages of HIS include enhancing signal coverage [21], improving communication rate [22], and boosting localization accuracy [23, 13]. However, due to its continuous nature, HIS incurs significant challenges for array signal processing, making traditional beamforming techniques [3, 4, 5, 6, 7, 8], which are tailored for discrete arrays, inefficient or even inapplicable [16, 17]. Due to these challenges, integrating HIS into the ISAC system, though promising for achieving a better performance trade-off between sensing and communication, is still in its infancy.
@@ I-A Prior work
Many efforts have been dedicated to beamforming design for the ISAC systems [3, 4, 5, 6, 24, 25, 26, 7, 8, 27, 28, 29], including transmit beamforming and joint transmit-receive beamforming. Transmit beamforming builds on the spatial degrees of freedom (DoFs) at the transmitter, which is devised to strike performance trade-off between sensing and communication [3, 4, 5, 24]. A commonly adopted strategy of transmit beamforming is to modify the transmit beam by optimizing the beamformer, aiming to maximize the power sent towards the targets while fulfilling the communication performance requirement at each user [3, 4, 5, 25, 24].
@@
Transmit beamforming fails to explore the spatial DoFs at the radar receiver, which holds potential to further enhance ISAC performance. Joint transmit-receive beamforming, in which the transmitter-side beamformers and radar receiver-side beamformers/filters are designed simultaneously, has been proposed in [27, 6, 7, 8, 26, 30, 28, 29]. The authors of [8] employed the Cramér-Rao bound (CRB) as the radar performance indicator and optimized the CRB while meeting the user signal-to-interference-and-noise ratio (SINR) requirement. However, explicit receive beamforming was not addressed. In [26], the authors considered transmit-receive beamforming in ISAC systems with SINR being the radar performance metric. However, without being optimized adaptively, the receive beamforming conducted at the radar receiver was solely determined by the steering vector. As a step further, [30] proposed a design framework for ISAC transmit-receive beamforming, optimizing transmit and receive beamformers alternately. The aforementioned works regarding ISAC beamforming typically rely on a discrete array, which limits performance due to inefficient utilization of a given aperture [14, 15].
@@
With the development of continuous-aperture array technology, HIS has been devised to alleviate the performance bottleneck caused by traditional discrete arrays, enabling higher array gain for a given aperture size [12, 13, 14, 15, 16, 17]. The superiority of HIS has been demonstrated in various systems aimed at different tasks [21, 22, 20, 18, 31, 32]. For instance, the channel model and performance bounds of HIS-based multi-user MIMO systems were analyzed in [33, 21, 34]. To characterize the radiation model of HIS, it is necessary to design the surface pattern, known as the current density distribution, for a continuous-aperture array [17, 16, 19, 18]. The authors of [35, 16, 17, 34, 18] modeled the radiation of continuous-aperture arrays and derived a closed-form expression of the capacity between two continuous volumes by utilizing the Fourier basis functions of continuous-space EM channels. To optimize the capacity of HIS-based communications, a Fourier transform-based pattern design approach was proposed by [19]. Specifically, a pattern division multiplexing (PDM) method was developed to model the radiation for HIS, aiming to maximize the sum rate when serving multiple communication users with HIS. Compared to [35, 16, 17, 34, 18] that employed the standard basis functions and projection coefficient matrices of the Fourier transform, the pattern design in [19] differed in incorporating additional optimizable projection coefficient matrices. This modification enhances the flexibility of HIS radiation modeling and pattern design, effectively addressing the issue of infinite computation resulting from the continuous nature of HIS.
@@ I-B Motivation and Contribution
Most existing works employing HIS have focused on communication tasks [19, 21, 22, 31, 32, 36], whereas the incorporation of HIS into ISAC systems has remained in its infancy [20, 37, 38]. Although the advantages of deploying sub-half-wavelength-spacing antennas for ISAC systems have been initially revealed in [20, 37], the advocated antenna array remained discrete. Therefore, the proposed schemes are not applicable to HISAC systems comprised of continuous-aperture arrays.
To fill the gap, this work focuses on incorporating HIS into ISAC system. Specifically, we propose an ISAC transceiver structure based on HIS, wherein both the transmitter and receiver of the base station (BS) are fabricated using HISs. By generating a specific pattern of the HIS, the radiation of the EM waves can be configured to meet the requirements of both sensing and communication. However, due to its continuity, the pattern of the HIS has infinite dimensions. Inspired by [19], we develop a continuous-discrete transformation based on the Fourier transform, which converts the continuous current from the space domain into the discrete wavenumber domain. Then, the problem of pattern design is transformed into a joint transmit-receive beamforming problem, where projection/beamforming matrices of the pattern in the wavenumber domain are optimizable to configure the beams of HISs. Based on this, we jointly design the transmit-receive beamformers by optimizing their coefficient matrices, maximizing the minimal sensing SINR among multiple targets while guaranteeing multi-user communication SINR requirements.
Due to the non-convex max-min objective function and coupled optimization variables of the joint transmit-receive design, the formulated problem is non-convex and, consequently, challenging to solve. We propose an alternating optimization (AO) algorithm to address it by designing the HIS transmit the receive beamforming iteratively. Specifically, we propose an adaptive bisection searching (ABS) method to accelerate the convergence of the semidefinite relaxation (SDR)-based bisection iteration process in the HIS transmit beamforming. Additionally, we propose a Rayleigh quotient-based algorithm for configuring the HIS receive beamforming. Numerical results demonstrate that our proposed HISAC system enhances the performance of both sensing and communication compared to the discrete array-based counterpart system.
@@
  • Continuous-Discrete Transformation: We convert the HIS pattern from the continuous space domain into the discrete wavenumber domain, through which the joint transmit-receive beamforming can be efficiently designed. Based on the continuous-discrete transformation, we formulate an optimization problem concerning the joint transmit-receive beamforming design for HISAC, balancing multi-target sensing performance while fulfilling multi-user communication performance requirements.
  • Joint Transmit-Receive Beamforming in HISAC: We develop an AO algorithm by recasting the original problem into: transmit HIS design and receive HIS configuration. Due to the high complexity of the involved SDR-based bisection search method, we propose an ABS method to obtain the refined potential range of the bisection search and accelerate the convergence in the iteration process. To address the receive HIS configuration, bluea Rayleigh quotient-based method is proposed obtaining a closed-form solution.
  • Performance: Numerical results reveal the significant performance gain of the proposed HISAC system, over its discrete-array-based counterpart. Specifically, the proposed schemes achieve a remarkable $9.5$ dB sensing SINR gain compared to the discrete-based arrays under the same communication SINR requirement.
@@ I-C Organization and Notation
The rest of the paper is organized as follows: Section II presents the HISAC system and formulates a pattern design problem. The continuous-discrete transformation is developed to convert the continuous pattern design into a joint transmit-receive beamforming problem in Section III. Section IV presents efficient methods to optimize patterns of both the transmit HIS and the receive HIS. Section V numerically demonstrates the solution and evaluates its performance in different settings. Finally, Section VI concludes the paper.
Throughout the paper, we use boldface lower-case and upper-case letters for vectors and matrices, respectively. The maximal eigenvalue of the matrix $\boldsymbol{A}$ is denoted by $\lambda_{max}(\boldsymbol{A})$. The $\ell_{2}$ norm, Frobenius norm, conjugate operation, transpose, Hermitian transpose, and stochastic expectation are written as $\|\cdot\|_{2}$, $\|\cdot\|_{F}$, $(\cdot)^{*}$, $(\cdot)^{T}$, $(\cdot)^{H}$, and $\mathbf{E}(\cdot)$, respectively. We use $\boldsymbol{I}_{N}$ to denote an $N$-dimensional identity matrix.
@@ II-A HISAC Transceiver Architecture
Figure 1: The geometry of the HISAC system.
Specifically, $\mathcal{P}_{S}$ sends signals to perform multi-target sensing and communicate to $K$ single-antenna users simultaneously, while $\mathcal{P}_{R}$ receives the radar echoes from $M$ targets. The structure of the HISAC transceiver operation is depicted in Fig. 2. The communication data streams and sensing streams are modulated into the transmit pattern at $\mathcal{P}_{S}$. Similarly, the echoed signals from $M$ targets are received by $\mathcal{P}_{R}$, and then demodulated by $M$ receive patterns.
@@
Let $\boldsymbol{x}=[\boldsymbol{c}^{T},\boldsymbol{s}^{T}]^{T}$, where $\boldsymbol{c}=[c_{1},..,c_{K}]^{T}\sim\mathcal{CN}\left(0,\boldsymbol{I}_{K}\right)$ and $\boldsymbol{s}=[s_{1},..,s_{M}]^{T}\sim\mathcal{CN}\left(0,\boldsymbol{I}_{M}\right)$ are communication data streams and sensing streams, respectively. Correspondingly, the transmit pattern at point $p$ of the surface is defined as $\boldsymbol{\rho}(p)=[\rho_{1},...,\rho_{K+M}]^{H}$. The pattern generated by the communication data streams is given by $j_{c}(p)=\sum_{k=1}^{K}c_{k}\rho_{k}(p)$ [19]. Similarly, the pattern generated by the sensing streams is given as $j_{r}(p)=\sum_{m=1}^{M}s_{m}\rho_{K+m}(p)$. The combined pattern at $\mathcal{P}_{S}$ is given by
@@
Similar to the transmit end, the received signal is filtered by a set of receive patterns $\boldsymbol{\tau}=[\tau_{1},...,\tau_{M}]^{H}$ at $\mathcal{P}_{R}$ to extract the target information.
The main purpose of the HISAC system design is to achieve maximal sensing performance while ensuring multiuser communication. To support multi-target sensing and multi-user communication, both the transmit pattern $\boldsymbol{\rho}$ and receive pattern $\boldsymbol{\tau}$ need to be optimized. We consider SINR as the performance metric of both communication and sensing, described as follows.

### v2

@@ Abstract
Traditional discrete-array-based systems fail to fully exploit interactions between closely spaced antennas, resulting in inadequate utilization of the aperture resource. In this paper, we propose a holographic intelligence surface (HIS) assisted integrated sensing and communication (HISAC) system, wherein both the transmitter and receiver are fabricated using a continuous-aperture array. To enable tractable design and optimization, we introduce a Fourier-based continuous-discrete transformation, converting the continuous pattern design into a finite-dimensional representation in the wavenumber domain. Based on this transformation, we formulate a joint transmit–receive beamforming optimization problem that aims to maximize multi-target sensing performance while satisfying multi-user communication requirements. To solve the non-convex problem with coupled variables, an alternating optimization-based algorithm is proposed to optimize the HISAC transmit-receive beamforming in an alternate manner. Specifically, the transmit beamforming design is solved by decoupling into a series of feasibility-checking sub-problems while the receive beamforming is determined by the Rayleigh quotient-based method. Simulation results demonstrate the superiority of the proposed HISAC system over traditional discrete-array-based ISAC systems, achieving significantly higher sensing performance while guaranteeing predetermined communication performance.
@@ I Introduction
isac system, which envisions simultaneous sensing and communication on a shared waveform/platform, is considered a vital solution for tackling challenges associated with spectral and/or power efficiency in 6G wireless networks [1, 2]. In a multi-antenna setting, joint ISAC waveform design is typically achieved through transmit beamforming design. This entails designing spatial beamformers judiciously to maximize sensing/communication under performance constraints in communication/sensing, thereby striking a balance between the two functionalities [3, 4, 5]. Furthermore, by leveraging radar receivers, joint transmit-receive beamforming has been proposed in the literature to further enhance ISAC performance [6, 7, 8].
The aforementioned beamforming designs in ISAC systems typically rely on discrete antenna arrays, where the antenna spacing is constrained to half-wavelength. This limits the number of antennas for a given aperture, thereby resulting in insufficient exploitation of array gain for ISAC performance enhancement [9, 10]. In addition, modeling electromagnetic (EM) propagation for the discrete array usually overlooks the mutual interference between two closely spaced antennas, potentially leading to inadequate information acquisition during signal reception [11, 12]. To overcome these limitations, the concept of holographic intelligence surface (HIS), also as known as holographic multiple-input multiple-output (MIMO) [13, 14] and continuous-aperture MIMO [15, 16], has emerged. By forming an effectively continuous-aperture, HIS enables more complete utilization of the physical aperture of the antenna array, thereby maximizing the radiated energy in the desired direction and approaching the aperture-limited directivity. Unlike the discrete array consisting of multiple patch antennas with half-wavelength spacing, HIS achieves a quasi-continuous-aperture array by deploying massive sub-wavelength tunable elements in a compact space [17]. This architecture mitigates the element-spacing constraints inherent in discrete arrays, enabling higher array gain for the same physical aperture size [10, 18]. The advantages of HIS include enhancing signal coverage [19], improving communication rate [20], and boosting localization accuracy [21, 10]. However, due to its continuous nature, HIS incurs significant challenges for array signal processing, making traditional beamforming techniques [3, 4, 5, 6, 7, 8], which are tailored for discrete arrays, inefficient or even inapplicable [13, 14]. Due to these challenges, integrating HIS into the ISAC system, though promising for achieving a better performance trade-off between sensing and communication, is still in its infancy.
@@ I-A Prior work
Many efforts have been dedicated to beamforming design for the ISAC systems [3, 4, 5, 6, 22, 7, 8, 23, 24], including transmit beamforming and joint transmit-receive beamforming. Transmit beamforming builds on the spatial degrees of freedom (DoFs) at the transmitter, which is devised to strike performance trade-off between sensing and communication [3, 4, 5]. A commonly adopted strategy of transmit beamforming is to modify the transmit beam by optimizing the beamformer, aiming to maximize the power sent towards the targets while fulfilling the communication performance requirement at each user [3, 4, 5].
@@
Transmit beamforming fails to explore the spatial DoFs at the radar receiver, which holds potential to further enhance ISAC performance. Joint transmit-receive beamforming, in which the transmitter-side beamformers and radar receiver-side beamformers/filters are designed simultaneously, has been proposed in [23, 6, 7, 8, 22, 25, 24]. In [22], the authors considered transmit-receive beamforming in ISAC systems with signal-to-interference-and-noise ratio (SINR) being the radar performance metric. However, without being optimized adaptively, the receive beamforming conducted at the radar receiver was solely determined by the steering vector. As a step further, [25] proposed a design framework for ISAC transmit-receive beamforming, optimizing transmit and receive beamformers alternately. The aforementioned works regarding ISAC beamforming typically rely on a discrete array, which limits performance due to inefficient utilization of a given aperture [11, 12].
TABLE I: Summary Comparison of Related HIS / Holographic Systems
Work Aperture Model System Scope Control Variables
Holographic MIMO [26] Discrete (element-level) Localization Phase shifts of RHS elements
Circular Holographic MIMO [27] Discrete (element-level) Comm. / Throughput Phase shifts of RHS elements
Overview of RHSs [28] Discrete (element-level) ISAC Amplitude and phase of RHS elements
Shape-adaptive RHSs [29] Discrete (element-level) Comm. / Throughput Amplitude, phase, and geometry parameters
RHS-assisted ISAC [17, 30] Discrete (element-level) ISAC Amplitude and phase of RHS elements
Continuous-aperture HIS [16, 31] Continuous Comm. / Throughput EM-consistent surface patterns
This work Continuous ISAC Wavenumber-domain coefficients
Note: Holographic MIMO, reconfigurable holographic surfaces (RHSs), and HISs denote related surface-based architectures. In particular, discrete-aperture HMIMO/RHS systems correspond to element-level control of amplitude and phase, whereas continuous-aperture HIS models rely on continuous electromagnetic representations or equivalent wavenumber-domain coefficients.
Abbreviations: Comm. : Communication.
With the development of continuous-aperture array technology, HIS has been devised to alleviate the performance bottleneck caused by traditional discrete arrays, enabling higher array gain for a given aperture size [9, 10, 11, 12, 13, 14]. The superiority of HIS has been demonstrated in various systems aimed at different tasks, including communication, localization, and other integrated services [32, 33, 26, 27, 28, 29, 17, 30]. Existing holographic systems are predominantly modeled using sub-wavenumber discrete dipole representations to approximate quasi-continuous distributions [28, 26, 27, 17, 30, 29]. Such formulations have been widely adopted for communication-centric designs [27, 26, 29] and ISAC systems [17, 30], as summarized in Table I. In particular, [29] adopts a discrete element-level model and focuses on optimizing surface geometry and phase shifts for communication-oriented performance. By contrast, continuous-aperture HIS models describe the surface using continuous electromagnetic fields, which provide a more fundamental characterization of the spatial DoF.
@@
To characterize the radiation model of HIS, it is necessary to design the surface pattern, known as the current density distribution, for a continuous-aperture array [14, 13, 16, 15]. The authors of [34, 35] modeled the radiation of continuous-aperture arrays and derived a closed-form expression of the capacity between two continuous volumes by utilizing the Fourier basis functions of continuous-space EM channels. To optimize the capacity of HIS-based communications, a pattern division multiplexing (PDM) method was developed to model the radiation for HIS [16], aiming to maximize the sum rate of multi-users with HIS.
@@ I-B Motivation and Contribution
Table I provides a high-level comparison of representative discrete-array, holographic, and HIS-based systems. Most existing works employing HIS have focused on communication tasks [31], whereas the incorporation of HIS into ISAC systems has remained in its infancy [17, 30]. Although the advantages of deploying sub-half-wavelength-spacing antennas for ISAC systems have been initially revealed in [17, 30], the advocated antenna array remained discrete. Therefore, the proposed schemes are not applicable to HISAC systems comprised of continuous-aperture arrays.
To fill the gap, this work focuses on incorporating HIS into ISAC systems. Specifically, we propose an ISAC transceiver structure based on HIS, in which both the transmitter and receiver at the base station (BS) are implemented using continuous-aperture HISs. By generating a specific pattern of the HIS, the radiation of the EM waves can be configured to support multi-target sensing and multi-user communication simultaneously. However, due to its continuity, the pattern of the HIS has infinite dimensions, posing significant computational challenges. To address this, we develop a continuous-discrete transformation based on the Fourier series, which maps the spatial-domain current into a finite-dimensional representation in the wavenumber domain. Compared to [16, 34], we further derive a closed-form expression for the far-field Green’s function in the Fourier domain, which enables precise SINR evaluation. This allows the original pattern design to be reformulated as a joint transmit–receive beamforming, where the beamforming matrices in the wavenumber domain are optimized to configure the radiation beams of the HIS. We then jointly optimize the transmit-receive beamformers to maximize the minimum sensing SINR across multiple radar targets, while ensuring that communication SINR requirements for all users are satisfied.
@@
  • Fourier-Domain Modeling of Continuous HIS Apertures: We propose a physically consistent Fourier-domain transformation that maps the continuous surface current distribution of a HIS into a finite-dimensional representation in the wavenumber domain. This leads to an equivalent steering vector expression for continuous-apertures, enabling direct alignment between HIS-based ISAC systems and conventional discrete-array models.
  • Discrete-Compatible Reformulation of the HISAC Problem: Leveraging the equivalent representation, we reformulate the originally infinite-dimensional beamforming problem into a finite-dimensional structure. The resulting optimization problem shares a similar mathematical form to those for discrete-array ISAC systems, which allows the direct application of existing optimization techniques. To support this formulation, we derive a closed-form expression for the Fourier transform of the far-field Green’s function, facilitating precise SINR evaluation in the HISAC problem.
  • Joint Transmit-Receive Beamforming in HISAC: We develop an alternating optimization (AO) algorithm by recasting the original problem into: transmit HIS design and receive HIS configuration. Due to the high complexity of the involved semidefinite relaxation (SDR)-based method, we propose an adaptive bisection searching (ABS) method to accelerate the convergence in the iteration process. To address the receive HIS configuration, a Rayleigh quotient-based method is proposed to obtain a closed-form solution.
@@ I-C Organization and Notation
The rest of the paper is organized as follows: Section II introduces the HISAC system and formulates the pattern design problem. Section III presents the continuous-discrete transformation and reformulates the problem as a joint transmit–receive beamforming design. Section IV describes the proposed optimization methods, and Section V provides numerical results. Finally, Section VI concludes the paper.
Throughout the paper, bold lowercase and uppercase letters denote vectors and matrices, respectively. Notations include $\lambda_{\max}(\boldsymbol{A})$ for the largest eigenvalue, $\|\cdot\|_{2}$ and $\|\cdot\|_{F}$ for the $\ell_{2}$ and Frobenius norms, $(\cdot)^{*}$, $(\cdot)^{T}$, and $(\cdot)^{H}$ for complex conjugate, transpose, and Hermitian, respectively, and $\mathbf{E}(\cdot)$ for expectation. The identity matrix of size $N$ is denoted by $\boldsymbol{I}_{N}$.
@@ II-A HISAC Transceiver Architecture
Figure 1: The illustration of the HIS-assisted integrated sensing and communication system.
Specifically, $\mathcal{P}_{S}$ sends signals to perform multi-target sensing and communicate to $K$ single-antenna users simultaneously, while $\mathcal{P}_{R}$ receives the radar echoes from $M$ targets. The structure of the HISAC transceiver operation is depicted in Fig. 2. The proposed HIS incorporates a single feed directly connected to the radio frequency (RF) distribution network within the surface. The traveling wave originates from the input port, propagates outward along the feed path, and excites the corresponding radiated wave [36]. By carefully designing the surface current distribution $j(p)$, the radiation characteristics $\psi_{\text{rad}}(p)$ can be flexibly controlled, enabling the desired EM environment for ISAC applications. Motivated by this capability, we propose to modulate both the communication data and sensing streams into the surface current pattern of the HIS. Specifically, the communication data streams and sensing streams are modulated into the transmit pattern at $\mathcal{P}_{S}$. Similarly, the echoed signals from $M$ targets are received by $\mathcal{P}_{R}$, and then demodulated by $M$ receive patterns.
@@
Let $\boldsymbol{x}=[\boldsymbol{c}^{T},\boldsymbol{s}^{T}]^{T}$, where $\boldsymbol{c}=[c_{1},..,c_{K}]^{T}\sim\mathcal{CN}\left(0,\boldsymbol{I}_{K}\right)$ and $\boldsymbol{s}=[s_{1},..,s_{M}]^{T}$ are communication data streams and sensing streams, respectively. In particular, the sensing signal sequence $\boldsymbol{s}$ is assumed to have unit power and be uncorrelated with other streams, which is generated by random phase coding [3]. Correspondingly, the transmit pattern at point $p$ of the surface is defined as $\boldsymbol{\rho}(p)=[\rho_{1},...,\rho_{K+M}]^{H}$. The pattern generated by the communication data streams is given by $j_{c}(p)=\sum_{k=1}^{K}c_{k}\rho_{k}(p)$ [16]. Similarly, the pattern generated by the sensing streams is given as $j_{r}(p)=\sum_{m=1}^{M}s_{m}\rho_{K+m}(p)$. The combined pattern at $\mathcal{P}_{S}$ is given by
@@
As shown in Fig. 2, by exciting each current cell at $\mathcal{P}_{S}$ with the modulated current $j(p)=\boldsymbol{\rho}^{H}(p)\boldsymbol{x}$, the ISAC signal is embedded into the HIS and subsequently radiated.
Similar to the transmit end, the received signal is filtered by a set of receive patterns $\boldsymbol{\tau}=[\tau_{1},...,\tau_{M}]^{H}$ at $\mathcal{P}_{R}$ to extract the target information. Each receive current cell $j(p^{\prime})=\boldsymbol{\tau}^{H}(p^{\prime})\boldsymbol{s}$ is designed to match the specific target, enabling accurate separation and generating the receive beams.
To fully exploit the reconfigurability of the HIS and maximize system performance, both $\boldsymbol{\rho}$ and $\boldsymbol{\tau}$ are jointly optimized to simultaneously improve sensing accuracy and increase communication capacity. In the following, we consider SINR as the performance metric of both communication and sensing, described as follows.